OpenStudy (anonymous):

Algebra Word Problem ( Work): A woman can do a certain job in 10 hours and her youngest daugther can do the job in 12 hours. After the woman and her youngest daugther work together for one hour, they are joined by the oldest daugther and the three of them finish the job in 3 more hours. How long does it take the oldest daugther to complete the job by herself?

OpenStudy (anonymous):

this one is tough. =)

OpenStudy (anonymous):

I know right :P

OpenStudy (ash2326):

@Maniqueen123 How much work the women will do in 1 hour? and her youngest daughter?

OpenStudy (anonymous):

you mean 1/10+1/12, but is that in one day...?

OpenStudy (anonymous):

actually no i thing the women does 1/10 in a hour and the youngest daughter does 1/12 in an hour

OpenStudy (ash2326):

That is the work done by them in an hour Women=1/10 youngest daughter=1/12 Let their middle daughter do=1/x work in an hour

OpenStudy (ash2326):

*her

OpenStudy (anonymous):

no you mean oldest, there is no middle

OpenStudy (ash2326):

Yeah:) Sorry

OpenStudy (ash2326):

Now try to create the equation, given that the women and youngest daughter work for 1 hour and then they are joined by the oldest daughter, and all three work for three hours and finish the work

OpenStudy (anonymous):

so is it like this mayb: 1/10 Woman, 1 hr. 1/12 Daugther, 1 hr. Together= 22/120

OpenStudy (ash2326):

Women and young daughter together \[\frac{1}{10}+\frac{1}{12}\] and then joined by the oldest and they work for three hours \[(\frac{1}{10}+\frac{1}{12}+\frac{1}{x})\times 3\] So total work= \[\frac{1}{10}+\frac{1}{12}+(\frac{1}{10}+\frac{1}{12}+\frac{1}{x})\times 3=???\] I've given you the hint, not tell me what this should be equal to

OpenStudy (anonymous):

22/120+(3/10+3/12+3/x)*3

OpenStudy (anonymous):

is that the first step...:/

OpenStudy (ash2326):

Yeah but what is it equal to??

OpenStudy (anonymous):

88/120=-3/x wait no...idk ...:(

OpenStudy (ash2326):

We have one work here, so it'd equal 1 \[\large \frac{1}{10}+\frac{1}{12}+(\frac{1}{10}+\frac{1}{12}+\frac{1}{x})\times 3=1\] Now you can find x?

OpenStudy (anonymous):

x=45/4